The SNeRG Bibliography

The SNePS Research Group Stuart C. Shapiro, Director William J. Rapaport, Associate Director Department of Computer Science 226 Bell Hall State University of New York at Buffalo Buffalo, NY 14260. snerg@cs.buffalo.edu

December 5, 1994

Copies of Departmental Technical Reports (marked with an asterisk (∗)) should be requested from The Library Committee, Department of Computer Science, SUNY at Buffalo, 226 Bell Hall, Buffalo, NY 14260. Copies of papers other than Departmental Technical Reports may be requested from Prof. Shapiro or Prof. Rapaport at the above address. Warning: Not all items are in chronological order.

References

[1] S. C. Shapiro. A net structure for semantic information storage, deduction and retrieval. In Proceedings of the Second International Joint Conference on Artificial Intelligence, pages 512–523, Los Altos, CA, 1971. Morgan Kaufmann.

[2] S. C. Shapiro. Generation as parsing from a network into a linear string. American Journal of Computational Linguistics, pages 45–62, 1975.

[3] S. C. Shapiro. An introduction to SNePS (semantic network processing system). Technical Report 31, Computer Science Department, Indiana University, Bloomington, IN, 1976.

[4] S. C. Shapiro and M. Wand. The relevance of relevance. Technical Report 46, Computer Science Department, Indiana University, Bloomington, IN, 1976.

[5] R. Bechtel and S. C. Shapiro. A logic for semantic networks. Technical Report 47, Computer Science Department, Indiana University, Bloomington, IN, 1976.

[6] S. C. Shapiro. Representing and locating deduction rules in a semantic network. Proceedings of the Workshop on Pattern-Directed Inference Systems, Sigart Newsletter(63):14–18, 1977.

[7] S. C. Shapiro. Representing numbers in semantic networks: prolegomena. In Proceedings of the 5th International Joint Conference on Artificial Intelligence, page 284, Los Altos, CA, 1977. Morgan Kaufmann.

[8] S. C. Shapiro. Compiling deduction rules from a semantic network into a set of processes. In Abstracts of Workshop on Automatic Deduction, page 7, Cambridge, MA, 1977.

[9] S. C. Shapiro. Path-based and node-based inference in semantic networks. In D. Waltz, editor, Tinlap– 2: Theoretical Issues in Natural Languages Processing, pages 219–225, New York, 1978. ACM.

[10] S. C. Shapiro. The SNePS semantic network processing system. In N. V. Findler, editor, Associative Networks: The Representation and Use of Knowledge by Computers, pages 179–203. Academic Press, New York, 1979.

[11] S. C. Shapiro. Generalized augmented transition network grammars for generation from semantic networks. In Proceedings of the 17th Annual Meeting of the Association for Computational Linguistics, pages 25–29. University of California at San Diego, 1979. Superseded by [34].

[12] S. C. Shapiro. Numerical quantifiers and their use in reasoning with negative information. In Proceedings of the Sixth International Joint Conference on Artificial Intelligence, pages 791–796, Los Altos, CA, 1979. Morgan Kaufmann.

[13] S. C. Shapiro. Using non-standard connectives and quantifiers for representing deduction rules in a semantic network, 1979. Invited paper presented at Current Aspects of AI Research, a seminar held at the Electrotechnical Laboratory, Tokyo.

[14] * D. P. McKay and S. C. Shapiro. MULTI: A LISP based multiprocessing system. Technical Report 164, Department of Computer Science, SUNY at Buffalo, 1980. (Contains appendices not in item 15).

[15] D. P. McKay and S. C. Shapiro. MULTI — a LISP based multiprocessing system. In Proceedings of the 1980 LISP Conference, pages 29–37. Stanford University, Stanford, CA, 1980.

[16] S. C. Shapiro and D. P. McKay. Inference with recursive rules. In ProceedIngs of the First Annual National Conference on Artificial Intelligence, pages 151–153, Los Altos, CA, 1980. Morgan Kaufmann.

[17] S. C. Shapiro. Review of Fahlman, Scott NETL: A system for representing and using real-world knowledge. American Journal of Computational Linguistics, 6(3):183–186, 1980.

[18] D. P. McKay. Recursive rules — an outside challenge. SNeRG Technical Note 1, Department of Computer Science, SUNY at Buffalo, 1980.

[19] * A. S. Maida and S. C. Shapiro. Intensional concepts in propositional semantic networks. Technical Report 171, Department of Computer Science, SUNY at Buffalo, 1981. Superseded by [41].

[20] * S. C. Shapiro. COCCI: A deductive semantic network program for solving microbiology unknowns. Technical Report 173, Department of Computer Science, SUNY at Buffalo, 1981.

[21] * J. Martins, D. P. McKay, and S. C. Shapiro. Bi-directional inference. Technical Report 174, Department of Computer Science, SUNY at Buffalo, 1981. Superseded by [40].

[22] * J. Martins and S. C. Shapiro. A belief revision system based on relevance logic and heterarchical contexts. Technical Report 175, Department of Computer Science, SUNY at Buffalo, 1981.

[23] S. C. Shapiro. Summary of scientific progress. SNeRG Technical Note 3, Department of Computer Science, SUNY at Buffalo, 1981.

[24] D. P. Mckay and J. Martins. SNePSLOG User’s Manual. Department of Computer Science, SUNY at Buffalo, 1981.

[25] D. P. McKay, H. Shubin, , and J. Martins. RIPOFF: Another text formatting program. SNeRG Technical Note 5, Department of Computer Science, SUNY at Buffalo, 1981. Obsolete.

[26] * J. Neal. A knowledge engineering approach to natural language understanding. Technical Report 179, Department of Computer Science, SUNY at Buffalo, 1981.

[27] * R. Srihari. Combining path-based and node-based reasoning in SNePS. Technical Report 183, Department of Computer Science, SUNY at Buffalo, 1981.

[28] D. P. McKay, J. Martins, E. Morgado, M. Almeida, and S. C. Shapiro. An assessment of SNePS for the Navy domain. SNeRG Technical Note 6, Department of Computer Science, SUNY at Buffalo, 1981.

[29] S. C. Shapiro. What do semantic network nodes represent? SNeRG Technical Note 7, Department of Computer Science, SUNY at Buffalo, 1981. Superseded by [41] and [101].

[30] D. P. McKay and S. C. Shapiro. Using active connection graphs for reasoning with recursive rules. In Proceedings of the Seventh International Joint Conference on Artificial Intelligence, pages 368–374, Los Altos, CA, 1981. Morgan Kaufmann.

[31] S. C. Shapiro and The SNePS Implementation Group. SNePS User’s Manual. Department of Computer Science, SUNY at Buffalo, 1981. Superseded by [207].

[32] S. C. Shapiro, D. P. McKay, J. Martins, and E. Morgado. SNePSLOG: A “higher order” logic programming language. SNeRG Technical Note 8, Department of Computer Science, SUNY at Buffalo, 1981. Presented at the Workshop on Logic Programming for Intelligent Systems, R.M.S. Queen Mary, Long Beach, CA.

[33] * H. Shubin. Inference and control in multiprocessing environments. Technical Report 186, Department of Computer Science, SUNY at Buffalo, 1981.

[34] S. C. Shapiro. Generalized augmented transition network grammars for generation from semantic networks. The American Journal of Computational Linguistics, 8(1):12–25, 1982.

[35] M. J. Almeida. NETP2 — a parser for a subset of english. SNeRG Technical Note 9, Department of Computer Science, SUNY at Buffalo, 1982.

[36] * L. M. Tranchell. A SNePS implementation of KL-ONE. Technical Report 198, Department of Computer Science, SUNY at Buffalo, 1982.

[37] S. C. Shapiro and J. G. Neal. A knowledge engineering approach to natural language understanding. In Proceedings of the 20th Annual Meeting of the Association for Computational Linguistics, pages 136–144, Menlo Park, CA, 1982. ACL.

[38] G. Donlon. Using resource limited inference in SNePS. SNeRG Technical Note 10, Department of Computer Science, SUNY at Buffalo, 1982.

[39] * J. T. Nutter. Defaults revisited or “Tell me if you’re guessing”. In Proceedings of the Fourth Annual Conference of the Cognitive Science Society, pages 67–69, Ann Arbor, MI, 1982. the Program in Cognitive Science of The University of Chicago and The University of Michigan.

[40] S. C. Shapiro, J. Martins, and D. McKay. Bi-directional inference. In Proceedings of the Fourth Annual Conference of the Cognitive Science Society, pages 90–93, Ann Arbor, MI, 1982. the Program in Cognitive Science of The University of Chicago and The University of Michigan.

[41] A. S. Maida and S. C. Shapiro. Intensional concepts in propositional semantic networks. Cognitive Science, 6(4):291–330, 1982. Reprinted in R. J. Brachman and H. J. Levesque, eds. Readings in Knowledge Representation, Morgan Kaufmann, Los Altos, CA, 1985, 170–189.

[42] J. P. Martins. Belief revision in MBR. In Proceedings of the 1983 Conference on Artificial Intelligence, Rochester, Michigan, 1983.

[43] * J. T. Nutter. What else is wrong with non-monotonic logics?: Representational and informational shortcomings. In Proceedings of the Fifth Annual Meeting of the Cognitive Science Society, page 5, Rochester, NY, 1983.

[44] M. J. Almeida and S. C. Shapiro. Reasoning about the temporal structure of narrative texts. In Proceedings of the Fifth Annual Meeting of the Cognitive Science Society, page 5, Rochester, NY, 1983.

[45] * J. P. Martins. Reasoning in Multiple Belief Spaces. PhD thesis, Technical Report 203, Department of Computer Science, SUNY at Buffalo, 1983.

[46] J. P. Martins and S. C. Shapiro. Reasoning in multiple belief spaces. In Proceedings of the Eighth International Joint Conference on Artificial Intelligence, pages 370–373, Los Altos, CA, 1983. Morgan Kaufmann.

[47] * J. T. Nutter. Default reasoning using monotonic logic: a modest proposal. In Proceedings of The National Conference on Artificial Intelligence, pages 297–300, Los Altos, CA, 1983. Morgan Kaufmann.

[48] * J. T. Nutter. Default reasoning in A.I. systems. Master’s thesis, Technical Report 204, Department of Computer Science, SUNY at Buffalo, 1983.

[49] E. J. Morgado and S. C. Shapiro. Believing and acting: An approach to meta-knowledge and metareasoning. SNeRG Technical Note 11, Department of Computer Science, SUNY at Buffalo, 1983. Superseded by [59].

[50] W. J. Rapaport and S. C. Shapiro. Quasi-indexical reference in propositional semantic networks. In Proceedings of Coling–84, pages 65–70. The Association for Computational Linguistics, 1984.

[51] * J. G. Neal and S. C. Shapiro. Knowledge-based parsing. Technical Report 213, Department of Computer Science, SUNY at Buffalo, 1984. Superseded by [98].

[52] * W. J. Rapaport. Belief representation and quasi-indicators. Master’s thesis, Technical Report 215, Department of Computer Science, SUNY at Buffalo, 1984. Superseded by [83].

[53] J. P. Martins and S. C. Shapiro. A model for belief revision. In Non-Monotonic Reasoning Workshop, pages 241–294. The American Association for Artificial Intelligence, 1984.

[54] Z. Xiang, S. N. Srihari, S. C. Shapiro, and J. G. Chutkow. Analogical and propositional representation of structure in neurological diagnosis. In Proc. First Conference on Artificial Intelligence Applications, pages 127–132, Silver Spring, MD, 1984. IEEE Computer Society Press.

[55] * J. G. Neal. A Knowledge Based Approach to Natural Language Understanding. PhD thesis, Technical Report 85–06, Department of Computer Science, SUNY at Buffalo, 1985.

[56] W. J. Rapaport. Meinongian semantics for propositional semantic networks. In Proceedings of the 23rd Annual Meeting of the Association for Computational Linguistics, pages 43–48, Morristown, NJ, 1985. Association for Computational Linguistics.

[57] V. Saks. A matcher for intensional semantic networks. SNeRG Technical Note 12, Department of Computer Science, SUNY at Buffalo, 1985.

[58] J. Suchin. A semantic network representation of the peripheral nervous system. SNeRG Technical Note 13, Department of Computer Science, SUNY at Buffalo, 1985.

[59] E. J. Morgado and S. C. Shapiro. Believing and acting: A study of meta-knowledge and meta-reasoning. In Proceedings of EPIA-85 (“Encontro Portugues de Inteligencia Artificial”), pages 138–154, Oporto, Portugal, 1985.

[60] Z. Xiang, S. N. Srihari, S. C. Shapiro, and J. G. Chutkow. A modeling scheme for diagnosis. In Expert Systems in Government Symposium, pages 538–547, Washington, DC, 1985. IEEE Computer Society Press.

[61] * S. C. Shapiro and W. J. Rapaport. SNePS considered as a fully intensional semantic network. Technical Report 85–15, Department of Computer Science, SUNY at Buffalo, 1985. Superseded by [101].

[62] * W. J. Rapaport. Logic and artificial intelligence. Technical Report 85–16, Department of Computer Science, SUNY at Buffalo, 1985.

[63] J. G. Neal and S. C. Shapiro. Parsing as a form of inference in a multiprocessing environment. In Proceedings of the Conference on Intelligent Systems and Machines, pages 19–24, Rochester, Michigan, 1985.

[64] Z. Xiang, J. G. Chutkow, S. C. Shapiro, and S. N. Srihari. Representation of spatial structure and function in diagnosis. In Proceedings of the Second Conference on Artificial Intelligence Applications, pages 223–228, Silver Spring, MD, 1985. IEEE Computer Society Press.

[65] *** M. Holynski, B. R. Gardner, and R. Ostrovsky. Towards an intelligent computer graphics system. Technical Report 86–003, Computer Science Department, Boston University, 1986.

[66] S. C. Shapiro, S. N. Srihari, J. Geller, and M.-R. Taie. A fault diagnosis system based on an integrated knowledge base. IEEE Software, 3(2):48–49, 1986.

[67] Z. Xiang, J. G. Chutkow, S. C. Shapiro, and S. N. Srihari. Computerized neurological diagnosis: A paradigm of modeling and reasoning. Health Care Instrumentation, 1(3):90–105, 1986.

[68] J. P. Martins and S. C. Shapiro. Theoretical foundations for belief revision. In J. Y. Halpern, editor, Theoretical Aspects of Reasoning About Knowledge, pages 383–398. Morgan Kaufmann Publishers, Los Altos, CA, 1986.

[69] S. C. Shapiro, S. N. Srihari, M.-R. Taie, and J. Geller. VMES: A network-based versatile maintenance expert system. In Applications of Artificial Intelligence to Engineering Problems: Proceedings of The 1st International Conference, pages 925–936, Berlin, 1986. Springer–Verlag.

[70] J. P. Martins and S. C. Shapiro. Hypothetical reasoning. In Applications of Artificial Intelligence to Engineering Problems: Proceedings of The 1st International Conference, pages 1029–1042, Berlin, 1986. Springer–Verlag.

[71] R. G. Hull. A new design for SNIP the SNePS inference package. SNeRG Technical Note 14, Department of Computer Science, SUNY at Buffalo, 1986.

[72] M. R. Taie, S. N. Srihari, J. Geller, and S. C. Shapiro. Device representation using instantiation rules and structural templates. In Proceedings of the Sixth Canadian Conference on Artificial Intelligence, pages 124–128. Presses de l’Université du Québec, 1986.

[73] J. P. Martins and S. C. Shapiro. Belief revision in SNePS. In Proceedings of the Sixth Canadian Conference on Artificial Intelligence, pages 230–234. Presses de l’Université du Québec, 1986.

[74] S. S. Campbell and S. C. Shapiro. Using belief revision to detect faults in circuits. SNeRG Technical Note 15, Department of Computer Science, SUNY at Buffalo, 1986.

[75] * W. J. Rapaport. Syntactic semantics: Foundations of computational natural-language understanding. Technical Report 86–24, Department of Computer Science, SUNY at Buffalo, 1986. Superseded by [103].

[76] * W. J. Rapaport, S. C. Shapiro, and J. M. Wiebe. Quasi-indicators, knowledge reports, and discourse. Technical Report 86–15, Department of Computer Science, SUNY at Buffalo, 1986.

[77] S. L. Hardt and W. J. Rapaport (eds.). Recent and current artificial intelligence research in the Department of Computer Science, SUNY Buffalo. AI Magazine, 7(2):91–100, 1986.

[78] S. C. Shapiro and W. J. Rapaport. SNePS considered as a fully intensional propositional semantic network. In Proc. Fifth National Conference on Artificial Intelligence, pages 278–283, Los Altos, CA, 1986. Morgan Kaufmann. Superseded by [101].

[79] * Ernesto J. M. Morgado. Semantic Networks as Abstract Data Types. PhD thesis, Technical Report 86-19, Department of Computer Science, State University of New York at Buffalo, Buffalo, NY, 1986. 318 pages.

[80] * G. A. Bruder, J. F. Duchan, W. J. Rapaport, E. M. Segal, S. C. Shapiro, and D. A. Zubin. Deictic centers in narrative: An interdisciplinary cognitive-science project. Technical Report 86–20, Department of Computer Science, SUNY at Buffalo, 1986. 32 pages.

[81] S. C. Shapiro. Symmetric relations, intensional individuals, and variable binding. Proceedings of the IEEE, 74(10):1354–1363, 1986.

[82] J. M. Wiebe and W. J. Rapaport. Representing de re and de dicto belief reports in discourse and narrative. Proceedings of the IEEE, 74(10):1405–1413, 1986.

[83] W. J. Rapaport. Logical foundations for belief representation. Cognitive Science, 10:371–422, 1986.

[84] S. C. Shapiro and J. Geller. Artificial intelligence and automated design. In The Computability of Design: Proceedings of the 1986 SUNY Buffalo Symposium on CAD, 1986. unpaginated, 13 pp.

[85] J. Van Blargan. An introduction to SNePS II. SNeRG Technical Note 16, Department of Computer Science, SUNY at Buffalo, 1986.

[86] C. M. Chan. Forward path-based inference in SNePS. SNeRG Technical Note 17, Department of Computer Science, SUNY at Buffalo, 1987.

[87] M. R. Taie, J. Geller, S. N. Srihari, and S. C. Shapiro. Knowledge based modeling of circuit boards. In Proceedings of the Annual Reliability and Maintainability Symposium, pages 422–427. IEEE, 1987.

[88] Naicong Li. Pronoun resolution in SNePS. SNeRG Technical Note 18, Department of Computer Science, SUNY at Buffalo, January 1986.

[89] * S. N. Srihari, W. J. Rapaport, and D. Kumar. On knowledge representation using semantic networks and sanskrit. Technical Report 87–03, Department of Computer Science, SUNY at Buffalo, 1987.

[90] J. G. Neal and S. C. Shapiro. Knowledge representation for reasoning about language. In J. C. Boudreaux, B. W. Hamill, and R. Jernigan, editors, The Role of Language in Problem Solving 2, pages 27–46. Elsevier Science Publishers, 1987.

[91] S. A. Chun. SNePS implementation of possessive phrases. SNeRG Technical Note 19, Department of Computer Science, SUNY at Buffalo, 1987.

[92] * A. H. Yuhan and S. C. Shapiro. Design of an incremental compiler for a production-system ATN machine. Technical Report 87–09, Department of Computer Science, SUNY at Buffalo, 1987.

[93] * M. J. Almeida. Reasoning About the Temporal Structure of Narratives. PhD thesis, Technical Report 87–10, Department of Computer Science, SUNY at Buffalo, 1987.

[94] ** J. T. Nutter. Assimilation: A strategy for implementing self-reorganizing knowledge bases. In Proceedings of the Sixth National Conference on Artificial Intelligence, pages 449–453, Los Angeles, CA, 1987. Morgan Kaufmann.

[95] S. L. Peters and S. C. Shapiro. A representation for natural category systems. In Proceedings of the Ninth Annual Conference of the Cognitive Science Society, pages 379–390. Lawrence Erlbaum Publishers, 1987.

[96] J. Geller and S. C. Shapiro. Graphical deep knowledge for intelligent machine drafting. In Proceedings of the Tenth International Joint Conference on Artificial Intelligence, pages 545–551, Los Angeles, CA, 1987. Morgan Kaufmann.

[97] S. L. Peters and S. C. Shapiro. A representation for natural category systems. In Proceedings of the Tenth International Joint Conference on Artificial Intelligence, pages 140–146, Los Altos, CA, 1987. Morgan Kaufmann.

[98] J. G. Neal and S. C. Shapiro. Knowledge-based parsing. In L. Bolc, editor, Natural Language Parsing Systems, pages 49–92. Springer–Verlag, Berlin, 1987.

[99] S. C. Shapiro and J. W. Rapaport. Knowledge representation for natural language processing. In 1987 Natural Language Planning Workshop, pages 56–77. Northeast Artificial Intelligence Consortium, Minnowbrook, NY, 1987. Superseded by [155].

[100] S. C. Shapiro and J. Geller. Artificial intelligence and automated design. In Y. Kalay, editor, Computability of Design, pages 173–187. John Wiley, 1987.

[101] S. C. Shapiro and W. J. Rapaport. SNePS considered as a fully intensional propositional semantic network. In N. Cercone and G. McCalla, editors, The Knowledge Frontier, pages 263–315. Springer– Verlag, New York, 1987.

[102] ** E. A. Fox, J. T. Nutter, T. Ahlswede, M. Evens, and J. Markowitz. Building a large thesaurus for information retrieval. In Proceedings of the Second Conference on Applied Natural Language Processing, pages 101–108, Austin, TX, 1988. ACL.

[103] W. J. Rapaport. Syntactic semantics: Foundations of computational natural-language understanding. In J. Fetzer, editor, Aspects of Artificial Intelligence, pages 81–131. Kluwer Academic Publishers, Dordrecht, Holland, 1988. Reprinted in E. Dietrich, editor, Thinking Computers & Virtual Persons: Essays on the Intentionality of Machines, pages 225–273. Academic Press, San Diego, 1994.

[104] S. C. Shapiro and W. J. Rapaport. Models and minds: A reply to Barnden. Northeast Artificial Intelligence Consortium Technical Report TR–8737, Department of Computer Science, SUNY at Buffalo, 1988. Superseded by [155].

[105] J. G. Neal and S. C. Shapiro. Intelligent multi-media interface technology. In J.W Sullivan and S.W. Tyler, editors, Proc. Architectures for Intelligent Interfaces: Elements and Prototypes, pages 69–91, Lockheed AI Center, 1988. Superseded by [147].

[106] J. P. Martins and S. C. Shapiro. A model for belief revision. Artificial Intelligence, 35(1):25–79, 1988.

[107] J. M. Wiebe and W. J. Rapaport. A computational theory of perspective and reference in narrative. In Proceedings of the 26th Annual Meeting of the Association for Computational Linguistics, pages 131–138. Association for Computational Linguistics, 1988.

[108] W. J. Rapaport. A knowledge-representation challenge for SNePS. SNeRG Technical Note 20, Department of Computer Science, SUNY at Buffalo, 1988.

[109] J. G. Neal, K. E. Bettinger, J. S. Byoun, Z. Dobes, and C. Y. Thielman. An intelligent multi-media human-computer dialogue system. In Proceedings of the Workshop on Space, Operations, Automation, and Robotics, (SOAR88), page 7. Wright State University, Dayton, OH, 1988.

[110] * L. D. Roberts and W. J. Rapaport. Quantifier order, reflexive pronouns, and quasi-indexicals: Essay, comments, and reply. Technical Report 88–16, Department of Computer Science, SUNY at Buffalo, 1988.

[111] S. L. Peters, S. C. Shapiro, and W. J. Rapaport. Flexible natural language processing and Roschian category theory. In Proceedings of the Tenth Annual Conference of the Cognitive Science Society, pages 125–131. Lawrence Erlbaum Publishers, 1988.

[112] J. G. Neal, Z. Dobes, K. E. Bettinger, and J. S. Byoun. Media-modual references in human-computer dialogue. In Proceedings of the Seventh National Conference of the American Association for Artificial Intelligence, pages 819–823. Morgan Kaufmann, 1988.

[113] S. C. Shapiro. Representing plans and acts. In Proceedings of the Third Annual Workshop on Conceptual Graphs, pages 3.2.7–1 – 3.2.7–6. The American Association for Artificial Intelligence, Menlo Park, CA, 1988.

[114] Deepak Kumar, Syed Ali, and Stuart C. Shapiro. Discussing, using and recognizing plans in SNePS preliminary report—SNACTor: An acting system. In P V S Rao and P Sadanandan, editors, Modern Trends in Information Technology: Proceedings of the Seventh Biennial Convention of South East Asia Regional Computer Confederation, pages 177–182. Tata McGraw-Hill, New Delhi, India, 1988.

[115] W. J. Rapaport. To think or not to think. Noûs, 22(4):585–609, 1988.

[116] S. C. Shapiro and The SNePS Implementation Group. SNePS-2 User’s Manual. Department of Computer Science, SUNY at Buffalo, 1989. Superseded by [207].

[117] S. C. Shapiro. Formal foundations of an intensional propositional semantic network. Presented at the Workshop on Formal Aspects of Semantic Networks, Santa Catalina Island, CA. Superseded by [148]., 1989.

[118] * W. J. Rapaport, E. M. Segal, S. C. Shapiro, D. A. Zubin, G. A. Bruder, J. F. Duchan, M. J. Almeida, J. H. Daniels, M. Galbraith, J. M. Wiebe, and A. H. Yuhan. Deictic centers and the cognitive structure of narrative comprehension. Technical Report 89–01, Department of Computer Science, SUNY at Buffalo, 1989.

[119] S. C. Shapiro, D. Kumar, and S. Ali. A propositional network approach to plans and plan recognition. In Proceedings of the 1988 Workshop on Plan Recognition, page 21, Los Altos, CA, 1989. Morgan Kaufmann.

[120] J. G. Neal, C. Y. Thielman, D. J. Funke, and J. S. Byoun. Multi-modal output composition for humancomputer dialogues. In Proceedings of the 1989 IEEE AI Systems in Government Conference, pages 250–257, George Washington Univ., Wash. D.C., 1989. IEEE.

[121] * W. J. Rapaport, E. M. Segal, S. C. Shapiro, D. A. Zubin, G. A. Bruder, J. F. Duchan, and D. M. Mark. Cognitive and computer systems for understanding narrative text. Technical Report 89–07, Department of Computer Science, SUNY at Buffalo, 1989.

[122] S. C. Shapiro. The CASSIE projects: An approach to natural language competence. In Proceedings of the 4th Portugese Conference on Artificial Intelligence, pages 362–380, Lisbon, Portugal, 1989. Springer-Verlag.

[123] J. G. Neal, C. Y. Thielman, Z. Dobes, S. M. Haller, and S. C. Shapiro. Natural language with integrated deictic and graphic gestures. In Proceedings of the DARPA Speech and Natural Language Workshop, page 14. Morgan Kaufmann, 1989.

[124] J. G. Neal, C. Y. Thielman, Z. Dobes, S. M. Haller, S. Glanowski, and S. C. Shapiro. CUBRICON: A multi-modal user interface. Presented at the GIS/LIS ’89 Conference, 1989.

[125] S. C. Shapiro and J. P. Martins. Recent advances and developments – the SNePS 2.1 report. In D. Kumar, editor, Current Trends in SNePS–Semantic Network Processing System: Proceedings of the First Annual SNePS Workshop, pages 1–13, Buffalo, NY, 1990. Springer-Verlag.

[126] M. R. Cravo and J. P. Martins. Path-based inference revisited. In D. Kumar, editor, Current Trends in SNePS–Semantic Network Processing System: Proceedings of the First Annual SNePS Workshop, pages 15–26, Buffalo, NY, 1990. Springer-Verlag.

[127] N. J. Mamede and J. P. Martins. Expanding SNePS capabilities with LORE. In D. Kumar, editor, Current Trends in SNePS–Semantic Network Processing System: Proceedings of the First Annual SNePS Workshop, pages 27–39, Buffalo, NY, 1990. Springer-Verlag.

[128] J. Geller. Order dependence of declarative knowledge representation. In D. Kumar, editor, Current Trends in SNePS–Semantic Network Processing System: Proceedings of the First Annual SNePS Workshop, pages 41–54, Buffalo, NY, 1990. Springer-Verlag.

[129] D. Kumar. An integrated model of acting and inference. In D. Kumar, editor, Current Trends in SNePS–Semantic Network Processing System: Proceedings of the First Annual SNePS Workshop, pages 55–65, Buffalo, NY, 1990. Springer-Verlag.

[130] R. R. Dipert. The structure of agency: Issues in the representation of agency and action. In D. Kumar, editor, Current Trends in SNePS–Semantic Network Processing System: Proceedings of the First Annual SNePS Workshop, pages 67–84, Buffalo, NY, 1990. Springer-Verlag.

[131] R. K. Srihari and W. J. Rapaport. Exacting visual information from text: Using captions to label human faces in newspaper photographs. In D. Kumar, editor, Current Trends in SNePS–Semantic Network Processing System: Proceedings of the First Annual SNePS Workshop, pages 85–96, Buffalo, NY, 1990. Springer-Verlag.

[132] J. T. Nutter. Knowledge based lexicons. In D. Kumar, editor, Current Trends in SNePS–Semantic Network Processing System: Proceedings of the First Annual SNePS Workshop, pages 97–106, Buffalo, NY, 1990. Springer-Verlag.

[133] W. J. Rapaport. Representing fiction in SNePS. In D. Kumar, editor, Current Trends in SNePS– Semantic Network Processing System: Proceedings of the First Annual SNePS Workshop, pages 107– 121, Buffalo, NY, 1990. Springer-Verlag.

[134] R. Wyatt. Kinds of opacity and their representations. In D. Kumar, editor, Current Trends in SNePS– Semantic Network Processing System: Proceedings of the First Annual SNePS Workshop, pages 123– 144, Buffalo, NY, 1990. Springer-Verlag.

[135] C. Rapp and M. Evens. Design of an emotion profiler using SNePS. In D. Kumar, editor, Current Trends in SNePS–Semantic Network Processing System: Proceedings of the First Annual SNePS Workshop, pages 145–152, Buffalo, NY, 1990. Springer-Verlag.

[136] B. E. Cline and J. T. Nutter. Implications of natural catagories for natural language generation. In D. Kumar, editor, Current Trends in SNePS–Semantic Network Processing System: Proceedings of the First Annual SNePS Workshop, pages 153–162, Buffalo, NY, 1990. Springer-Verlag.

[137] D. Kumar, S. S. Ali, J. Haas, and S. C. Shapiro. The SNePS acting system. In K. E. Bettinger and G. Srikantan, editors, Proceedings of the Fifth Annual University at Buffalo Graduate Conference on Computer Science, pages 91–100, 1990.

[138] * J. M. Wiebe. Recognizing Subjective Sentences: A Computational Investigation of Narrative Text. PhD thesis, Technical Report 90-03, Department of Computer Science, SUNY at Buffalo, 1990.

[139] * S. C. Shapiro and W. J. Rapaport. Models and minds: Knowledge representation for naturallanguage competence. Technical Report 90–10, Department of Computer Science, SUNY at Buffalo, 1990. Superseded by [155].

[140] * William J. Rapaport. Predication, fiction, and artificial intelligence. Technical Report 90–11, Department of Computer Science, SUNY at Buffalo, 1990. Superseded by [146].

[141] J. M. Lammens. Cultural literacy: Educating Cassie? SNeRG Technical Note 21, Department of Computer Science, SUNY at Buffalo, 1990.

[142] Susan M. Haller and Syed S. Ali. Using focus for generating felicitous locative expressions. In Proceedings of the Third International Conference on Industrial and Engineering Applications of Artificial Intelligence, page 6, Charleston, SC, 1990. Association for Computing Machinery. To appear.

[143] * S. C. Shapiro and W. J. Rapaport. The SNePS family. Technical Report 90–21, Department of Computer Science, SUNY at Buffalo, 1990. Superseded by [157].

[144] * S. C. Shapiro. Case studies of SNePS. Technical Report 90–29, Department of Computer Science, SUNY at Buffalo, 1990. Superseded by [149].

[145] Sandra L. Peters and William J. Rapaport. Superordinate and basic level categories in discourse: Memory and context. In Program of the Twelfth Annual Conference of the Cognitive Science Society, pages 157–165, Hillsdale, NJ, 1990. Lawrence Erlbaum Associates.

[146] William J. Rapaport. Predication, fiction, and artificial intelligence. Topoi, 10:79–111, 1991.

[147] Jeannette G. Neal and Stuart C. Shapiro. Intelligent multi-media interface technology. In Joseph W. Sullivan and Sherman W. Tyler, editors, Intelligent User Interfaces, pages 11–43. Addison Wesley, Reading, MA, 1991.

[148] Stuart C. Shapiro. Cables, paths and “subconsious” reasoning in propositional semantic networks. In John F. Sowa, editor, Principles of Semantic Networks, chapter 4, pages 137–156. Morgan Kaufmann, San Mateo, CA, 1991.

[149] Stuart C. Shapiro. Case studies of SNePS. SIGART Bulletin, 2(3):128–134, June 1991.

[150] * Stuart C. Shapiro, Hans Chalupsky, and Hsueh-Cheng Chou. Connecting ARC/INFO and SNACTor. Technical Report 91-13, Department of Computer Science, SUNY at Buffalo, Buffalo, NY, July 1991. Superseded by [164].

[151] Deepak Kumar and Stuart C. Shapiro. Architecture of an intelligent agent in SNePS. SIGART Bulletin, 2(4):89–92, August 1991.

[152] Christopher M. Lusardi. A SNePS/Cassie implementation of parts of the dictionary of cultural literacy. SNeRG Technical Note 22, Department of Computer Science, SUNY at Buffalo, Buffalo, NY, September 1991. 74 pages.

[153] Deepak Kumar and Stuart C. Shapiro. Modeling a rational cognitive agent in SNePS. In P. Barahona, L. Moniz Pereira, and A. Porto, editors, EPIA 91: 5th Portugese Conference on Artificial Intelligence, Lecture Notes in Artificial Intelligence 541, pages 120–134. Springer-Verlag, Heidelberg, 1991.

[154] Joongmin Choi and Stuart C. Shapiro. Experience-based deductive learning. In Third International Conference on Tools for Artificial Intelligence TAI ’91, pages 502–503. IEEE Computer Society Press, Los Alamitos, CA, 1991.

[155] Stuart C. Shapiro and William J. Rapaport. Models and minds: Knowledge representation for naturallanguage competence. In Robert Cummins and John Pollock, editors, Philosophy and AI: Essays at the Interface, pages 215–259. MIT Press, Cambridge, MA, 1991.

[156] Joongmin Choi and Stuart C. Shapiro. Efficient implementation of non-standard connectives and quantifiers in deductive reasoning systems. In Proceedings of the Twenty-Fifth Hawaii International Conference on System Sciences, pages 381–390. IEEE Computer Society Press, Los Alamitos, CA, 1992.

[157] Stuart C. Shapiro and William J. Rapaport. The SNePS family. Computers & Mathematics with Applications, 23(2–5):243–275, January–March 1992.

[158] Stuart C. Shapiro and William J. Rapaport. SNePS considered as a fully intensional propositional semantic network. In Leslie Burkholder, editor, Philosophy and the Computer, pages 75–91. Westview Press, Boulder, CO, 1992. Slightly revised version of [78]. Superseded by [101].

[159] Stuart C. Shapiro. Relevance logic in computer science. Section 83 of Alan Ross Anderson and Nuel D. Belnap, Jr. and J. Michael Dunn et al. Entailment, Volume II, pages 553–563. Princeton University Press, Princeton, NJ, 1992.

[160] * Henry Hexmoor, Johan Lammens, and Stuart C. Shapiro. An autonomous agent architecture for integrating perception and acting with grounded, embodied symbolic reasoning. Technical Report 92-21, Department of Computer Science, SUNY at Buffalo, Buffalo, NY, September 1992. 22 pages.

[161] Martin J. Zaidel. XGinseng: An X windows graphic display tool for SNePS networks. SNeRG Technical Note 23, Department of Computer Science, SUNY at Buffalo, Buffalo, NY, May 1991. 18 pages.

[162] Daniel Campos. Temporal relations and spanish narrative: A small step towards language understanding. SNeRG Technical Note 24, Department of Computer Science, SUNY at Buffalo, Buffalo, NY, September 1991. 136 pages.

[163] John S. Lewocz. XGinseng: An X windows editing and display tool for SNePS networks. SNeRG Technical Note 25, Department of Computer Science, SUNY at Buffalo, Buffalo, NY, August 1992. 9 pages.

[164] Stuart C. Shapiro, Hans Chalupsky, Hsueh-Cheng Chou, and David M. Mark. Intelligent user interfaces: Connecting ARC/INFO and SNACTor, a semantic network based system for planning actions. In Proceedings of the Twelfth Annual ESRI User Conference, V. 3, pages 151–165. Environmental Systems Research Institute, Redlands, California, 1992.

[165] * Henry Hexmoor and Donald Nute. Methods for deciding what to do next and learning. Technical Report 92-23, Department of Computer Science, SUNY at Buffalo, Buffalo, NY, September 1992. 20 pages.

[166] Deepak Kumar. A unified model of acting and inference. In Proceedings of the Twenty-Sixth Hawaii International Conference on System Sciences. IEEE Computer Society Press, Los Alamitos, CA, 1993.

[167] Deepak Kumar. An AI architecture based on message passing. In James Geller, editor, Proceedings of The 1993 AAAI Spring Symposium on Innovative Applications of Massively Parallel Architectures, pages 127–131. AAAI Press, March 1993.

[168] Deepak Kumar. Rational engines for BDI architectures. In Amy Lansky, editor, Proceedings of The 1993 AAAI Spring Symposium on Foundations of Automated Planning, pages 78–82. AAAI Press, March 1993.

[169] Henry Hexmoor, Guido Caicedo, Frank Bidwell, and Stuart C. Shapiro. Air battle simulation: An agent with ‘conscious’ and ‘unconscious’ layers. In Daniel F. Boyd, editor, UBGCCS 93: Proceedings of The Eighth Annual University at Buffalo Graduate Conference on Computer Science, Technical Report No. 93-14, pages 52–59. Department of Computer Science, State University of New York at Buffalo, Buffalo, NY, March 1993.

[170] Karen Ehrlich and William J. Rapaport. Vocabulary expansion through natural-language context. In Daniel F. Boyd, editor, UBGCCS 93: Proceedings of The Eighth Annual University at Buffalo Graduate Conference on Computer Science, Technical Report No. 93-14, pages 78–84. Department of Computer Science, State University of New York at Buffalo, Buffalo, NY, March 1993.

[171] Henry Hexmoor, Johan Lammens, and Stuart C. Shapiro. Embodiment in GLAIR: a grounded layered architecture with integrated reasoning for autonomous agents. In Douglas D. Dankel II and John Stewman, editors, Proceedings of The Sixth Florida AI Research Symposium (FLAIRS 93), pages 325–329. the Florida AI Research Society, April 1993.

[172] * Johan Lammens, Henry Hexmoor, and Stuart C. Shapiro. Of elephants and men. Technical Report 93-13, Department of Computer Science, SUNY at Buffalo, Buffalo, NY, April 1993. 26 pages.

[173] * Henry Hexmoor, Johan Lammens, Guido Caicedo, and Stuart C. Shapiro. Behavior based ai, cognitive processes, and emergent behaviors in autonomous agents. Technical Report 93-15, Department of Computer Science, SUNY at Buffalo, Buffalo, NY, April 1993. 15 pages.

[174] Susan M. Haller. Interactive generation of plan justifications. In Preprints of the Fourth European Workshop on Natural Language Generation, pages 79–89. CNR—Consiglio Nazionale delle Ricerche and AI*IA—Associazione Italiana per l’Intelligenza Artificiale, April 1993.

[175] Stuart C. Shapiro. Belief spaces as sets of propositions. Journal of Experimental and Theoretical Artificial Intelligence (JETAI), 5(2&3):225–235, April–September 1993.

[176] Deepak Kumar and Stuart C. Shapiro. Deductive efficiency, belief revision and acting. Journal of Experimental and Theoretical Artificial Intelligence (JETAI), 5(2&3):167–177, April–September 1993.

[177] Syed S. Ali. A structured representation for noun phrases and anaphora. In Proceedings of the Fifteenth Annual Conference of the Cognitive Science Society, pages 197–202, Hillsdale, NJ, June 1993. Lawrence Erlbaum.

[178] Deepak Kumar. The SNePS BDI architecture. Journal of Decision Support Systems—Special Issue on Logic Modeling, 1994. Forthcoming.

[179] Stuart C. Shapiro. Preface to special issue on knowledge representation for natural language processing. Minds and Machines, 3(4):377–380, November 1993.

[180] Syed S. Ali and Stuart C. Shapiro. Natural language processing using a propositional semantic network with structured variables. Minds and Machines, 3(4):421–451, November 1993.

[181] Susan M. Haller. Collaboration in an interactive model of plan explanation. In Proceedings of the AAAI Fall Symposium on Human-Computer Collaboration, Raliegh, North Carolina. Technical Report FS93-05, AAAI, October 1993.

[182] Susan M. Haller. Planning for intentions with rhetorical relations. In Owen Rambow, editor, Intentionality and Structure in Discourse Relations, pages 23–26. Association for Computational Linguistics, June 1993.

[183] Susan M. Haller. An interactive model for plan explanation. In Chris Rowles, Huan Liu, and Norman Foo, editors, AI ’93: Proceedings of the 6th Australian Joint Conference on Artificial Intelligence, pages 371–376. World Scientific, Singapore, November 1993.

[184] Hans Chalupsky. Using hypothetical reasoning as a method for belief ascription. Journal of Experimental and Theoretical Artificial Intelligence (JETAI), 5(2&3):119–133, April–September 1993.

[185] * Richard W. Wyatt. The representation of opaque contexts. Master’s thesis, Technical Report 89–13, Department of Computer Science, State University of New York at Buffalo, Buffalo, NY, December 1989.

[186] Susan M. Haller. A model for cooperative interactive plan explanation. In Proceedings of the The Tenth IEEE Conference on Artificial Intelligence for Applications, pages 254–260, Los Alamitos, CA, 1994. IEEE Computer Society Press.

[187] * James Geller. A Knowledge Representation Theory for Natural Language Graphics. PhD thesis, Technical Report 88-15, Department of Computer Science, State University of New York at Buffalo, Buffalo, NY, July 1988.

[188] * William J. Rapaport. Computer processes and virtual persons: Comments on Cole’s ‘Artificial intelligence and personal identity’. Technical Report 90-13, Department of Computer Science, State University of New York at Buffalo, Buffalo, NY, May 1990.

[189] * Albert Hanyong Yuhan. Dynamic Computation of Spatial Reference Frames in Narrative Understanding. PhD thesis, Technical Report 91-03, Department of Computer Science, State University of New York at Buffalo, Buffalo, NY, February 1991.

[190] * Rohini Srihari. Extracting Visual Information from Text: Using Captions to Label Faces in Newspaper Photographs. PhD thesis, Technical Report 91-17, Department of Computer Science, State University of New York at Buffalo, Buffalo, NY, November 1991.

[191] * Joongmin Choi. Experienced-Based Learning in Deductive Reasoning Systems. PhD thesis, Technical Report 93-20, Department of Computer Science, State University of New York at Buffalo, Buffalo, NY, May 1993.

[192] * Johan M. Lammens and Stuart C. Shapiro. Learning symbolic names for perceived colors. Technical Report 93-31, Department of Computer Science, State University of New York at Buffalo, Buffalo, NY, August 1993.

[193] * Henry H. Hexmoor, Johan Lammens, and Stuart C. Shapiro. An autonomous agent architecture for integrating “unconscious” and “conscious” reasoned behaviors. Technical Report 93-37, Department of Computer Science, State University of New York at Buffalo, Buffalo, NY, August 1993.

[194] Douglas J. Funke, Jeannette G. Neal, and Rajendra D. Paul. An approach to intelligent automated window management. Int. J. Man-Machine Studies, 38:949–983, 1993.

[195] * Syed S. Ali. A “Natural Logic” for Natural Language Processing and Knowledge Representation. PhD thesis, Technical Report 94-01, Department of Computer Science, State University of New York at Buffalo, Buffalo, NY, January 1994.

[196] * Deepak Kumar. From Beliefs and Goals to Intentions and Actions: An Amalgamated Model of Inference and Acting. PhD thesis, Technical Report 94-04, Department of Computer Science, State University of New York at Buffalo, Buffalo, NY, 1994.

[197] * Henry H. Hexmoor. What are routines good for? Technical Report 94-07, Department of Computer Science, State University of New York at Buffalo, Buffalo, NY, February 1994.

[198] * Henry H. Hexmoor. A methodology for developing competent agents without sensor and actuator profusion. Technical Report 94-09, Department of Computer Science, State University of New York at Buffalo, Buffalo, NY, March 1994.

[199] * Robin K. Hill. Issues of Semantics in a Semantic-Network Representation of Belief. PhD thesis, Technical Report 94-11, Department of Computer Science, State University of New York at Buffalo, Buffalo, NY, April 1994.

[200] Susan M. Haller. Representing discourse for efficient interactive generation. In Douglas D. Dankel II, editor, Proceedings of the Seventh Florida Artificial Intelligence Research Symposium, pages 179–183. the Florida AI Research Society, St. Petersburg, FL, May 1994.

[201] Deepak Kumar and Stuart C. Shapiro. Acting in service of inference (and vice versa). In Douglas D. Dankel II, editor, Proceedings of the Seventh Florida Artificial Intelligence Research Symposium, pages 207–211. the Florida AI Research Society, St. Petersburg, FL, May 1994.

[202] * Johan M. Lammens. A Computational Model of Color Perception and Color Naming. PhD thesis, Technical Report 94-26, Department of Computer Science, State University of New York at Buffalo, Buffalo, NY, June 1994.

[203] * William J. Rapaport. Understanding understanding: Syntactic semantics and computational cognition. Technical Report 94-28, Department of Computer Science, State University of New York at Buffalo, July 1994.

[204] Hans Chalupsky and Stuart C. Shapiro. SL: A subjective, intensional logic of belief. In Proceedings of the Sixteenth Annual Conference of the Cognitive Science Society, pages 165–170, Hillsdale, NJ, August 1994. Lawrence Erlbaum.

[205] Jeannette G. Neal and Stuart C. Shapiro. Knowledge-based multimedia systems. In John F. Koegel Buford, editor, Multimedia Systems, pages 403–438. ACM Press/Addison Wesley, Reading, MA, 1994.

[206] Henry Hexmoor, Johan Lammens, Guido Caicedo, and Stuart C. Shapiro. Behaviour based AI, cognitive processes, and emergent behaviors in autonomous agents. In G. Rzevski, J. Pastor, and R. Adey, editors, Applications of Artificial Intelligence in Engineering VIII, volume 2, pages 447–461. Computational Mechanics Publications with Elsevier Applied Science, Southampton and London, 1993.

[207] S. C. Shapiro and The SNePS Implementation Group. SNePS 2.1 User’s Manual. Department of Computer Science, SUNY at Buffalo, 1994.

[208] * Susan M. Haller. Interactive Generation of Plan Descriptions and Justifications. PhD thesis, Technical Report 94-40, Department of Computer Science, State University of New York at Buffalo, Buffalo, NY, November 1994.

[209] Johan M. Lammens, Henry H. Hexmoor, and Stuart C. Shapiro. Of elephants and men. In Luc Steels, editor, Proceedings of the NATO ASI on the Biology and Technology of Intelligent Autonomous Agents. Springer Verlag, in press.

File: /projects/snwiz/Bibliography/bibliography.dvi